# Bottleneck-Minimal Indexing for Generative Document Retrieval

### v1

@@ 1 Introduction
For such reconsideration, it is natural to apply the rate-distortion theory initiated by Shannon (Shannon, 1948, 1959; Berger, 1968), which is based on the mutual information. Let $X$, $T$, and $Q$ denote random variables defined over $\mathcal{X}$, $\mathcal{T}$, and $\mathcal{Q}$, respectively. As the rate-distortion theory only involves two terms corresponding to $X$ and $T$, the aim, in short, is to acquire $T$ that minimally distorts $X$. All previous GDR proposals thus consider its distortion optimality, as illustrated in Figure 1(b1).
@@ 3.2 Distortion Optimality
where $\beta$ is the Lagrange multiplier. This coefficient $\beta$ determines the relative importance of $I(T;Q)$ in minimizing $\mathcal{L}$. When $\beta\to 0$, the latter term can be ignored, and $T$ shrinks to a single point so that the former term decreases to 0. Conversely, as $\beta$ becomes large, $T$ tends to simply copy $X$. Because the possibilities for $X$ are observed via the subset $\mathcal{X}$, a large $\beta$ corresponds to observing more documents.
@@ 3.4 A Theoretical Solution to Optimality
$p(\mathcal{X},\mathcal{Q}|f)\equiv\prod_{x\in\mathcal{X}}p^{*}(T=f(x)|X=x),$ (6)
where $p^{*}(T|X)$ is given by Formula (5).
@@ 4.1 Hierarchical Random Indexing (HRI)
We consider the most basic method, random indexing. For each document, an ID digit is randomly selected according to some prior distribution $p(T)$. $\mathcal{X}$ is partitioned into $|V|$ subsets, where $m$, the length of $t$, is 1; this case is called atomic (Tay et al., 2022) if $|V|\geq|\mathcal{X}|$. Furthermore, each subset is recursively partitioned for representation by $V$. The recursive subdivision constitutes a hierarchy of subdivisions of depth $m$.
@@ 4.2 Hierarchical $k$-Means Indexing (HKmI)
Hence, $k$-means provides a BMI if we assume $p(Q|X)\sim N(\mu_{x},\Sigma)$ and $p(Q|T)\sim N(\mu_{t},\sigma I)$, where $N$ represents a Gaussian distribution; $\mu_{x}$ and $\mu_{t}$ are mean vectors specific to $x$ and $t$, respectively; $\Sigma$ is any covariance matrix; and $\sigma I$ is a diagonal matrix. This conclusion immediately follows from substituting the Gaussian density functions into Formula (5). The optimality holds for any choice of $\beta$ in Formula (3). A proof is provided in Appendix A.2. The linkage between $k$-means clustering and information bottleneck theory has been theoretically explored by Still et al. (2003) in depth.
@@ 4.4 Our Proposal: Bottleneck-Minimal Indexing (BMI)
In a real application, the distributions $p(Q|X=x)$ and $p(Q|T=t)$ may deviate from Gaussian assumptions, potentially exhibiting multi-modal characteristics. Under these conditions, $k$-means clustering methods become suboptimal and should be replaced by another sophisticated clustering methods that capture reflect the true distribution of the data. Developing such methods form a future direction.
@@ 5.2 Model and Training Settings
We used the same neural-network architecture as in NCI (Wang et al., 2022). For sequence-to-sequence generation of an index for a query, the NCI architecture combined a standard transformer (Vaswani et al., 2017) encoder with a prefix-aware weight-adaptive (PAWA) decoder. The encoder parameters were initialized with T5 weights (Raffel et al., 2020) acquired by large-scale pretraining, and the decoder weights were randomly initialized.
We tested models of different sizes, which were initialized from the weights of T5-tiny [^2] https://huggingface.co/google/t5-efficient-tiny, T5-mini [^3] https://huggingface.co/google/t5-efficient-mini, T5-small [^4] https://huggingface.co/t5-small, and T5-base [^5] https://huggingface.co/t5-base, where the string after “T5” indicates the weights. The PAWA decoder had two transformer layers for T5-tiny and T5-mini, three for T5-small, and six for T5-base. All models were trained using the default hyperparameters of NCI, as provided in its official GitHub repository666github.com/solidsea98/Neural-Corpus-Indexer-NCI.
@@ 6.1 Quantification of GDR Bottleneck
Next, Figure 4 compares the random (black) and $k$-means (blue) hierarchical indexing methods in terms of the Rec@1 on the NQ320K test set. As the model size increased, the Rec@1 gradually increased for both indexing methods. With increasing size, the performances of the random and $k$-means methods became competitive. For the T5-base model, the hierarchical random method even outperformed hierarchical $k$-means. This indicates that when T5-base is large enough, the model associate an index string directly with a document, without requiring semantic compression of the document space.
@@ 6.4 Comparison with SOTA
Figure 5 shows the results of the finetuning process. Solid curves shows the performance using enhanced GenQ queries, while dashed curves represent the baseline results reported in the second column of Table 2. The HKmI method was also examined for its use of enhanced GenQ queries as augmented training data, although it did not apply these for indexing as our method does.
@@
These findings underscore our method’s effectiveness and simplicity, utilizing static, number-based indexes generated through a simple $k$-means process. Our model’s training procedure is the same as that of NCI, without necessitating additional computational resources, with the sole distinction being the production of document indexes as by incorporating the query distribution.
@@ Appendix A Mathematical Details
A.2 $k$-Means Is BMI
Under the Gaussian assumption stated in Section 4.2, Formula (5) simplifies as follows:
$p^{*}(T|X)$ $=\frac{p^{*}(T)}{Z(X,\beta)}\exp\Bigl{(}-\beta\text{KL}\bigl{[}N(\mu_{x},\Sigma)\bigm{\lVert}N(\mu_{t},\sigma I)\bigr{]}\Bigr{)},$ (17)
$=\frac{p^{*}(T)}{Z(X,\beta)}\exp\Bigl{(}-\beta\text{KL}\bigl{[}N(\mu_{x},\Sigma)\bigm{\lVert}N(\mu_{t_{i}},\sigma I)\bigr{]}\Bigr{)},$ (18)
$=\frac{p^{*}(T)}{Z(X,\beta)}\exp\left\{-\frac{\beta}{2}\left[\log\frac{|\sigma I|}{|\Sigma|}-d+\text{tr}\bigl{(}(\sigma I)^{-1}\Sigma\bigr{)}+(\mu_{t}-\mu_{x})^{\text{T}}(\sigma I)^{-1}(\mu_{t}-\mu_{x})\right]\right\},$ (19)
$=\frac{p^{*}(T)}{Z(X,\beta)}\exp\left\{-\frac{\beta}{2\sigma}\lVert\mu_{t}-\mu_{x}\rVert^{2}\right\}\underbrace{\exp\left\{-\frac{\beta}{2}\Bigl{(}d\log\sigma-d-\log|\Sigma|+\text{tr}\bigl{(}\Sigma\bigr{)}/\sigma\Bigr{)}\right\},}_{\text{constant}}$ (20)
where $d$ is the vector space’s dimensionality.
In other words, a BMI indexing produces
$p(\mathcal{T}|\mathcal{X})$ $\propto\prod_{x\in\mathcal{X}}p(T)\exp\left\{-\frac{\beta}{2\sigma}\lVert\mu_{t}-\mu_{x}\rVert^{2}\right\}$ (21)
$=p(\mathcal{T})\exp\left\{-\frac{\beta}{2\sigma}\sum_{(x,t)\in(\mathcal{X},\mathcal{T})}\lVert\mu_{t}-\mu_{x}\rVert^{2}\right\}.$ (22)
@@ A.2 $k$-Means Is BMI
On the other hand, the $k$-means clustering method aims to minimize the sum of the squared distances from each cluster’s centroid to the cluster members. Let $g:x\mapsto i$ represent the partition of points into $k$ clusters, denoted as $[1,2,\cdots,k]$. Then, $k$-means clustering is formulated as follows (Hartigan & Wong, 1979):
$\min_{g}S(g)$ $=\sum_{i=1}^{k}\sum_{x\in g^{-1}(i)}\lVert x-\mu_{i}\rVert^{2},$ (23)
$\text{where}~{}\mu_{i}$ $=\frac{1}{|g^{-1}(i)|}\sum_{x\in g^{-1}(i)}x.$ (24)
This objective function is equivalent to the maximum-likelihood objective function in the following:
$\max_{g}p(\mathcal{X}|\mathcal{T})$ $=\prod_{x\in\mathcal{X}}N(\mu_{x}|\mu_{t},\sigma I)$ (25)
@@
$=\frac{1}{(2\pi\sigma)^{|\mathcal{X}|/2}}\exp\left(-\frac{1}{2\sigma^{2}}S(g)\right).$ (26)
Finally, by Bayes’ theorem, we have
@@
$p(\mathcal{T}|\mathcal{X})=\frac{p(\mathcal{X}|\mathcal{T})p(\mathcal{T})}{p(\mathcal{X})}=\frac{1}{(2\pi\sigma)^{|\mathcal{X}|/2}p(\mathcal{X})}p(\mathcal{T})\exp\left(-\frac{1}{2\sigma^{2}}S(g)\right),$ (27)
where corresponds to Formula (22) if we set $\mu_{x}$ to $x$ and $\mu_{t}$ to the centroid as calculated following Formula (24).

### v2

@@ 1 Introduction
For such reconsideration, it is natural to apply the rate-distortion theory initiated by Shannon (Shannon, 1948, 1959; Berger, 1968), which is based on the mutual information. Let $X$, $T$, and $Q$ denote random variables defined over $\mathcal{X}$, $\mathcal{T}$, and $\mathcal{Q}$, respectively. As the rate-distortion theory only involves two terms corresponding to $X$ and $T$, the aim, in short, is to acquire $T$ that minimally distorts $X$. All previous GDR proposals thus consider its distortion optimality, as illustrated in Figure 1(b-1).
@@ 3.2 Distortion Optimality
where $\beta$ is the Lagrange multiplier. This coefficient $\beta$ determines the relative importance of $I(T;Q)$ in $\mathcal{L}$. When $\beta\to 0$, the latter term can be ignored, and $T$ shrinks to a single point so that the former term decreases to 0. Conversely, as $\beta$ becomes large, $T$ tends to simply copy $X$. Because the possibilities for $X$ are observed via the subset $\mathcal{X}$, a large $\beta$ corresponds to observing more documents.
@@ 3.4 A Theoretical Solution to Optimality
$p(\mathcal{X},\mathcal{Q}|f)\equiv\prod_{x\in\mathcal{X}}p^{*}\bigl{(}X=x\bigm{|}T=f(x)\bigr{)}.$ (6)
where $p^{*}$ is given by Formula (5).
@@ 4.1 Hierarchical Random Indexing (HRI)
We consider the most basic method, random indexing. For each document, an ID digit is randomly selected according to some prior distribution $p(T)$. $\mathcal{X}$ is partitioned into $|V|$ subsets. Furthermore, each subset is recursively partitioned for representation by $V$. The recursive subdivision constitutes a hierarchy of subdivisions of depth $m$. A special case arises when $|V|\geq|\mathcal{X}|$ and $m=1$; such indexing is termed atomic (Tay et al., 2022), since $\mathcal{T}$ lacks hierarchical structure.
@@ 4.2 Hierarchical $k$-Means Indexing (HKmI)
Hence, $k$-means provides a BMI if we assume $p(Q|X)\sim N(\mu_{x},\Sigma)$ and $p(Q|T)\sim N(\mu_{t},\sigma^{2}I)$, where $N$ represents a Gaussian distribution; $\mu_{x}$ and $\mu_{t}$ are mean vectors specific to $x$ and $t$, respectively; $\Sigma$ is any covariance matrix; and $\sigma^{2}I$ is a diagonal matrix. This conclusion immediately follows from substituting the Gaussian density functions into Formula (5). The optimality holds for any choice of $\beta$ in Formula (3). A proof is provided in Appendix A.2. The linkage between $k$-means clustering and information bottleneck theory has been theoretically explored by Still et al. (2003) in depth.
@@ 4.4 Our Proposal: Bottleneck-Minimal Indexing (BMI)
In a real application, the distributions $p(Q|X=x)$ and $p(Q|T=t)$ may deviate from Gaussian assumptions, potentially exhibiting multi-modal characteristics. Under these conditions, $k$-means clustering methods become suboptimal and should be replaced by another sophisticated clustering methods that reflect the true distribution of the data. Developing such methods form a future direction.
@@ 5.2 Model and Training Settings
We used the same neural-network architecture as in NCI (Wang et al., 2022). For sequence-to-sequence generation of an index for a query, the NCI architecture combined a standard transformer (Vaswani et al., 2017) with a prefix-aware weight-adaptive (PAWA) decoder. The Transformer encoder’s parameters were initialized with T5 weights (Raffel et al., 2020) acquired by large-scale pretraining, while the decoder weights were randomly initialized.
We tested models of different sizes, which were initialized from the weights of T5-tiny [^2] https://huggingface.co/google/t5-efficient-tiny, T5-mini [^3] https://huggingface.co/google/t5-efficient-mini, T5-small [^4] https://huggingface.co/t5-small, and T5-base [^5] https://huggingface.co/t5-base, where the string after “T5” indicates the weights. The PAWA decoder had four transformer layers. All models were trained using the default hyperparameters of NCI, as provided in its official GitHub repository666github.com/solidsea98/Neural-Corpus-Indexer-NCI.
@@ 6.1 Quantification of GDR Bottleneck
Next, Figure 4 compares the random (black) and $k$-means (blue) hierarchical indexing methods in terms of the Rec@1 on the NQ320K test set. As the model size is reduced, Rec@1 shows a general decline for both methods. With the reduction in size, the superiority of the $k$-means method over the random method became apparent. While the performance of the two methods was similar for the T5-base model, the Rec@1 of the $k$-means method was more than twice that of the random method in the T5-tiny model.
@@ 6.4 Comparison with SOTA
Figure 5 shows the results of the finetuning process. Solid curves shows the performance using enhanced GenQ queries, while dashed curves represent the baseline results reported in the second column of Table 2. The HKmI method was also examined for its use of enhanced GenQ queries as augmented training data, although it, unlike our method, did not utilize these enhanced GenQ queries for indexing.
@@
These findings underscore our method’s effectiveness and simplicity, utilizing static, number-based indexes generated through a simple $k$-means process. Our model’s training procedure is the same as that of NCI, without necessitating additional computational resources, with the sole distinction being the production of document indexes by incorporating query distribution.
@@ Appendix A Mathematical Details
A.2 $k$-Means Produces Bottleneck Minimum
In both the HKmI (Section 4.2) and BMI (Section 4.4), the $k$-means clustering method is applied hierarchically to the vector representations of documents $\mathcal{X}$. We demonstrate that at each hierarchical level, as $\mathcal{X}$ is assigned identifiers from the alphabet $V$, $k$-means clustering effectively minimizes the information bottleneck as defined in Definition 6.
@@ A.2 $k$-Means Produces Bottleneck Minimum
Let $\mathbf{x}$ represent the vector representation of a document $x$. The goal of $k$-means clustering is to partition $\{\mathbf{x}:x\in\mathcal{X}\}$ into $k$ clusters, aiming to minimize the sum of the squared distances from each document to its cluster’s centroid. This is represented by the partition function $f:x\mapsto i$. The optimization problem is described as follows (Hartigan & Wong, 1979):
$\min_{f}S(f)$ $=\sum_{i=1}^{k}\sum_{x\in f^{-1}(i)}\lVert\mathbf{x}-\bar{\mathbf{x}}_{i}\rVert^{2},$ (17)
$\text{where}~{}\bar{\mathbf{x}}_{i}$ $=\frac{1}{|f^{-1}(i)|}\sum_{x\in f^{-1}(i)}\mathbf{x}.$ (18)
This paper focuses on the optimal solution $f^{*}$ of Formula (17). We acknowledge various implementations for achieving $f^{*}$; however, they are not covered in this paper.
We now aim to establish the following proposition:
Proposition A.3.
Assume that $p(Q|x)\sim N(\mathbf{x},\Sigma)$ and $p(Q|t)\sim N(\mu_{t},\sigma^{2}I)$, where $\Sigma$ is an arbitrary covariance matrix and $\sigma^{2}I$ is an isotropic diagonal matrix; $\mu_{t}$ is the vector estimated for every element $t\in V$ in the alphabet $V$, Then, any optimal solution $f^{*}$ to Formula (17) also maximizes the likelihood $p(\mathcal{X},\mathcal{Q}|f)$ as defined in Formula (6).
Proof.
By definition, $p(\mathcal{X},\mathcal{Q}|f)=\prod_{x\in\mathcal{X}}p^{*}(X=x\mid T=f(x))$, where $p^{*}$ is detailed in Formula (5). From Bayes’ theorem:
$p^{*}(X|T)=\frac{p^{*}(T|X)}{p^{*}(T)}p(X),$ (19)
where $p(X)$ is a constant independent of $f$.
Applying Formula (19) and then inserting Formula (5), we derive:
$p(\mathcal{X},\mathcal{Q}|f)$ $\propto\prod_{x\in\mathcal{X}}\frac{p^{*}(T=f(x)|X=x)}{p^{*}(T=f(x))}$ (20)
$=\frac{1}{\prod_{x\in\mathcal{X}}Z(x,\beta)}\exp\left(-\beta\sum_{x\in\mathcal{X}}\text{KL}\Bigl{[}p(Q|X=x)\Bigm{\lVert}p\bigl{(}Q|T=f(x)\bigr{)}\Bigr{]}\right)$ (21)
Assuming $p(Q|x)\sim N(\mathbf{x},\Sigma)$ and $p(Q|t)\sim N(\mu_{t},\sigma^{2}I)$, the KL divergence is computed analytically; expanding the expression in Formula (21) leads to:
$p(\mathcal{X},\mathcal{Q}|f)$ $=p(\mathcal{X},\mathcal{Q}|f,\{\mu_{t}:t\in V\})$ (22)
$\propto\exp\left\{-\frac{\beta}{2\sigma^{2}}\sum_{x\in\mathcal{X}}\lVert\mathbf{x}-\mu_{f(x)}\rVert^{2}\right\}\underbrace{\exp\left\{-\frac{\beta}{2}|\mathcal{X}|\Bigl{(}d\log\sigma^{2}-d-\log|\Sigma|+\text{tr}\bigl{(}\Sigma\bigr{)}/\sigma^{2}\Bigr{)}\right\},}_{\text{constant}}$ (23)
@@
$\propto\exp\left\{-\frac{\beta}{2\sigma^{2}}\sum_{x\in\mathcal{X}}\lVert\mathbf{x}-\mu_{f(x)}\rVert^{2}\right\}$ (24)
@@
$=\exp\left\{-\frac{\beta}{2\sigma^{2}}\sum_{t\in V}\sum_{x\in f^{-1}(t)}\lVert\mathbf{x}-\mu_{t}\rVert^{2}\right\}.$ (25)
Since $\beta$ and $\sigma$ are constants, maximizing $p(\mathcal{X},\mathcal{Q}|f)$ equates to minimizing $\sum_{t\in V}\sum_{x\in f^{-1}(t)}\lVert\mathbf{x}-\mu_{t}\rVert^{2}$. At the maximum, $\mu_{t}$ corresponds to the centroid $\bar{\mathbf{x}}_{t}$. This directly aligns with the $k$-means objective in Formula (17) and the selection of centroid as the cluster center in Formula (18).
Therefore, any optimal solution $f^{*}$ to Formula (17) also optimally maximizes the likelihood $p(\mathcal{X},\mathcal{Q}|f)$.
The effectiveness of $k$-means in HKmI or BMI is validated through Proposition A.3. In HKmI, assuming $p(Q|x)\sim N(\mu_{x},\Sigma)$, where $\mu_{x}$ represents the BERT vector representation for the document, aligns with maximizing $p(\mathcal{X},\mathcal{Q}|f)$. For BMI, where the assumption is $p(Q|x)\sim N(\mu_{Q|x},\Sigma)$, we also achieve maximization of $p(\mathcal{X},\mathcal{Q}|f)$.
However, it is a simplification in HKmI to assume $p(Q|x)\sim N(\mu_{x},\Sigma)$ as $\mu_{x}$ describes the document, not necessarily its query distribution. In contrast, BMI’s assumption of $p(Q|x)\sim N(\mu_{Q|x},\Sigma)$ is more suitable.